# Not All Clients Are Equal: Personalized Federated Learning on Heterogeneous Multi-Modal Clients

### v1

@@ A.6 Comparison of Memory and Computational Costs
FedMKT (COLING 2025) - 1.000 Logit Extraction 1.574
FedMosaic (Ours) PQ-LoRA 1.053 Last Layer Gradient Compute 1.098
@@
A.7 Empirical Analysis of Block-Wise Aggregation
@@
To identify layer-wise correspondences between depth-heterogeneous models, we analyze representation alignment using CKA (kornblith2019similarity, ). Specifically, we measure similarity across layers within the LLaMA-3 family (1B, 3B, 8B) and the Qwen-2.5 family (0.5B, 1.5B, 3B), as illustrated in Fig. 8. As shown in the figure, layers with the same relative depth exhibit strong alignment, indicating approximately linear alignment within both the LLaMA-3 and Qwen-2.5 families. Moreover, we observe near-linear alignment even across families, i.e., between LLaMA-3 and Qwen-2.5, despite weaker linearity than intra-family alignment. This empirical analysis supports our block-wise aggregation of PQ-LoRA. We provide an illustration of the block-wise PQ-LoRA in Fig. 9.
@@
A.8 Experimental Results in More Diverse Heterogeneous PFL Scenarios
@@
Method DRAKE-Dynamic DRAKE-Static
@@
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
SFT 66.41$\pm$0.84 59.17$\pm$0.64 47.94$\pm$0.18 47.10$\pm$0.19 68.86$\pm$2.16 62.30$\pm$3.12 47.98$\pm$0.27 47.44$\pm$0.20
DITTO (ICML 2021) 61.56$\pm$0.01 56.36$\pm$0.20 48.19$\pm$0.26 47.41$\pm$0.01 65.48$\pm$0.97 60.34$\pm$0.32 48.49$\pm$0.03 48.09$\pm$0.01
FedSim (ICML 2022) 65.00$\pm$0.41 58.18$\pm$0.46 47.47$\pm$0.10 46.61$\pm$0.04 67.39$\pm$0.32 62.42$\pm$0.25 47.08$\pm$0.05 47.10$\pm$0.03
FedIT (ICASSP 2024) 66.18$\pm$0.22 58.95$\pm$0.02 47.54$\pm$0.02 46.97$\pm$0.01 68.78$\pm$0.22 64.07$\pm$0.92 47.34$\pm$0.55 47.33$\pm$0.29
TAKFL (NeurIPS 2024) 64.73$\pm$0.49 58.33$\pm$0.53 47.85$\pm$0.46 47.16$\pm$0.08 68.17$\pm$0.75 63.19$\pm$1.21 47.28$\pm$0.14 46.90$\pm$1.00
FedDPA (NeurIPS 2024) 63.26$\pm$0.09 57.31$\pm$0.02 48.04$\pm$0.26 47.21$\pm$0.07 67.71$\pm$0.95 63.03$\pm$0.24 48.18$\pm$0.53 48.22$\pm$0.13
FedDAT (AAAI 2024) 60.43$\pm$0.54 56.82$\pm$0.29 49.79$\pm$0.21 48.25$\pm$0.25 61.97$\pm$0.58 57.64$\pm$0.45 49.76$\pm$0.46 47.99$\pm$0.21
PerAda (CVPR 2024) 58.10$\pm$2.74 54.84$\pm$1.68 46.90$\pm$0.22 46.69$\pm$0.21 61.57$\pm$0.97 56.22$\pm$0.38 45.01$\pm$0.37 47.06$\pm$0.71
FedMKT (COLING 2025) 63.25$\pm$0.05 57.23$\pm$0.18 47.88$\pm$0.29 47.25$\pm$0.15 65.50$\pm$0.37 61.12$\pm$0.56 48.27$\pm$0.04 47.70$\pm$0.31
FedMosaic (Ours) 68.94$\pm$0.68 60.96$\pm$0.06 52.18$\pm$0.34 50.11$\pm$0.03 70.41$\pm$1.27 65.12$\pm$1.15 52.67$\pm$0.20 50.91$\pm$0.70
@@
A.9 Experiment Results on Homogeneous PFL Setup
In addition to the heterogeneous PFL scenario (Sec. 6.2), where both data distributions and model architectures vary across clients, we also evaluate FedMosaic in a homogeneous PFL setting, where clients share the same model architecture but have heterogeneous data distributions. We assume all clients use LLaVA-3B models and summarize the results for the PFL-Dynamic scenario in Tab.13 and for the PFL scenario in Tab.13. Consistent with results in the heterogeneous setup, FedMosaic significantly outperforms baselines in generalization ability (i.e., ‘Others’) and even surpasses SFT in personalization performance (i.e., ‘Self’).
Method DRAKE-Dynamic HFLB-Dynamic
@@ A.9 Experiment Results on Homogeneous PFL Setup
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
SFT 67.57$\pm$0.14 59.18$\pm$1.26 49.70$\pm$0.02 48.04$\pm$1.18 80.26$\pm$1.14 77.98$\pm$0.49 63.24$\pm$0.23 62.90$\pm$0.60
DITTO (ICML 2021) 64.09$\pm$0.55 57.22$\pm$0.01 49.92$\pm$0.27 49.05$\pm$0.07 79.34$\pm$0.41 76.72$\pm$0.84 64.10$\pm$0.40 63.57$\pm$0.06
FedSim (ICML 2022) 67.05$\pm$0.98 60.69$\pm$0.84 47.94$\pm$1.69 47.55$\pm$1.14 79.60$\pm$0.77 76.93$\pm$0.78 60.85$\pm$1.15 60.50$\pm$0.82
FedIT (ICASSP 2024) 67.62$\pm$0.49 60.35$\pm$0.69 49.84$\pm$0.19 48.95$\pm$0.13 79.34$\pm$0.59 78.40$\pm$1.21 62.51$\pm$1.30 62.38$\pm$0.96
TAKFL (NeurIPS 2024) 67.02$\pm$0.04 59.86$\pm$0.36 50.03$\pm$0.06 48.85$\pm$0.09 79.45$\pm$0.69 77.15$\pm$0.99 63.01$\pm$0.66 62.70$\pm$0.96
FedDPA (NeurIPS 2024) 66.59$\pm$0.86 58.68$\pm$0.71 49.92$\pm$0.24 48.81$\pm$0.13 80.57$\pm$0.62 77.65$\pm$0.61 63.58$\pm$0.30 63.02$\pm$0.09
FedDAT (AAAI 2024) 61.95$\pm$0.90 57.22$\pm$0.45 51.50$\pm$0.13 50.28$\pm$0.08 79.58$\pm$0.53 76.82$\pm$1.13 67.33$\pm$0.13 66.47$\pm$0.29
PerAda (CVPR 2024) 63.78$\pm$0.70 57.44$\pm$0.16 49.79$\pm$0.25 49.10$\pm$0.02 78.88$\pm$0.96 76.31$\pm$1.24 63.35$\pm$1.26 63.09$\pm$0.56
FedMKT (COLING 2025) 65.20$\pm$0.43 58.73$\pm$0.12 49.68$\pm$0.29 48.84$\pm$0.22 79.43$\pm$0.69 77.24$\pm$0.61 62.66$\pm$0.38 62.79$\pm$0.86
FedMosaic (Ours) 70.80$\pm$0.23 62.18$\pm$0.64 53.93$\pm$0.36 51.72$\pm$0.28 80.75$\pm$0.14 78.87$\pm$0.20 68.30$\pm$0.69 67.46$\pm$0.22
Method DRAKE-Static HFLB-Static
@@
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
SFT 70.56$\pm$0.27 66.31$\pm$0.37 50.37$\pm$0.04 49.99$\pm$0.16 80.89$\pm$0.85 80.01$\pm$0.12 63.40$\pm$0.41 63.09$\pm$0.17
DITTO (ICML 2021) 66.19$\pm$0.38 62.05$\pm$0.69 50.59$\pm$0.05 49.95$\pm$0.08 80.37$\pm$0.37 79.05$\pm$0.25 63.93$\pm$0.24 63.52$\pm$0.10
FedSim (ICML 2022) 68.65$\pm$0.86 63.96$\pm$0.78 49.27$\pm$0.22 48.85$\pm$0.14 80.05$\pm$0.01 79.36$\pm$0.79 61.22$\pm$0.30 61.31$\pm$0.12
FedIT (ICASSP 2024) 70.14$\pm$0.03 66.08$\pm$0.59 50.09$\pm$0.15 49.87$\pm$0.26 80.63$\pm$1.23 79.97$\pm$0.01 63.34$\pm$0.11 63.09$\pm$0.04
TAKFL (NeurIPS 2024) 68.75$\pm$0.04 64.39$\pm$0.07 49.70$\pm$0.16 49.60$\pm$0.03 79.97$\pm$0.11 79.41$\pm$0.81 63.67$\pm$0.39 63.30$\pm$0.25
FedDPA (NeurIPS 2024) 67.40$\pm$2.73 63.71$\pm$1.24 50.45$\pm$0.38 49.92$\pm$0.30 79.99$\pm$1.23 79.02$\pm$0.35 63.45$\pm$0.07 63.11$\pm$0.02
FedDAT (AAAI 2024) 64.09$\pm$1.16 61.10$\pm$0.60 52.20$\pm$0.26 51.47$\pm$0.04 79.80$\pm$0.50 78.43$\pm$0.68 68.92$\pm$0.44 67.66$\pm$0.27
PerAda (CVPR 2024) 65.64$\pm$0.48 61.93$\pm$0.58 50.53$\pm$0.02 49.94$\pm$0.05 80.25$\pm$0.45 79.01$\pm$0.26 64.00$\pm$0.13 63.58$\pm$0.31
FedMKT (COLING 2025) 67.85$\pm$0.92 62.93$\pm$1.16 49.23$\pm$1.54 48.77$\pm$1.69 79.85$\pm$0.67 79.72$\pm$0.60 63.84$\pm$0.20 63.39$\pm$0.07
FedMosaic (Ours) 72.02$\pm$0.37 67.18$\pm$0.13 54.21$\pm$0.05 52.94$\pm$0.05 81.69$\pm$0.16 80.54$\pm$0.09 68.70$\pm$0.26 67.75$\pm$0.23
@@
A.10 Experiment Results on Qwen-based LLaVA
@@
As shown in the table, FedMosaic significantly outperforms baselines in both homogeneous and heterogeneous PFL scenarios, consistent with results from LLaVA-Llama3 variants. This demonstrates that our proposed PQ-LoRA consistently facilitates knowledge sharing across heterogeneous models and SIMA reduces interference during local model aggregation, regardless of architecture.
Method DRAKE-Homonegeous DRAKE-Heterogeneous
@@
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
SFT 68.79$\pm$1.36 59.06$\pm$0.18 47.87$\pm$0.40 47.10$\pm$0.04 65.79$\pm$0.17 58.24$\pm$0.14 44.40$\pm$0.86 44.06$\pm$0.43
FedIT (ICASSP 2024) 68.40$\pm$0.07 60.58$\pm$0.41 49.70$\pm$0.52 48.54$\pm$0.07 66.02$\pm$0.72 58.26$\pm$1.06 44.47$\pm$0.30 44.06$\pm$0.43
FedDPA (NeurIPS 2024) 66.73$\pm$0.42 60.21$\pm$0.47 49.39$\pm$0.07 48.69$\pm$0.11 65.43$\pm$0.64 57.79$\pm$0.08 44.48$\pm$0.43 44.53$\pm$0.26
FedMKT (COLING 2025) 65.95$\pm$0.35 58.81$\pm$0.64 49.73$\pm$0.09 48.52$\pm$0.24 62.85$\pm$0.99 56.71$\pm$0.33 44.73$\pm$1.35 44.24$\pm$0.75
FedMosaic (Ours) 70.38$\pm$0.34 62.84$\pm$0.23 54.25$\pm$0.33 51.85$\pm$0.20 67.36$\pm$0.21 59.95$\pm$0.71 50.43$\pm$0.01 48.51$\pm$0.29
@@
A.11 Experimental Results on Fed-LLM
In addition to Fed-Aya (Sec.6.2), we evaluate FedMosaic on another LLM-FL benchmark, Fed-LLM, and summarize the results in Tab. 14. As shown in the table, FedMosaic outperforms baselines in both personalization and generalization, consistent with results from other MLLM-FL benchmarks, such as DRAKE and HFLB.
Method Self Others
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
SFT 25.21$\pm$1.31 29.35$\pm$0.61 25.83$\pm$1.03 27.30$\pm$0.73
FedIT (ICASSP 2024) 24.06$\pm$0.83 28.42$\pm$0.45 26.67$\pm$0.66 28.28$\pm$0.38
FedDPA (NeurIPS 2024) 24.93$\pm$1.06 27.51$\pm$0.55 26.72$\pm$0.67 27.69$\pm$0.27
FedMKT (COLING 2025) 24.92$\pm$0.96 27.51$\pm$0.63 29.55$\pm$0.43 30.57$\pm$0.19
FedMosaic (Ours) 31.24$\pm$0.13 33.84$\pm$0.53 29.26$\pm$0.33 31.70$\pm$0.19
Table 14: Quantitative comparison of heterogeneous LLM clients on the FedLLM-Static. Three clients use the LLaMA-3B model, while the remaining two use the LLaMA-8B model.
A.12 Client-Wise Accuracy

### v3

@@ A.8 Comparison of Computational and Memory Costs
FedMKT (COLING 2025) Public data Logit Share 1.002 Logit Extraction 1.574
FedMosaic (Ours) PQ-LoRA 1.053 Last Layer Gradient Compute & Client-wise Similarity Compute 1.098
@@
We compare the computational cost $\mathcal{C}$ and memory cost $\mathcal{M}$ of various baselines and summarize the results in Tab. 15. Following (Seo et al., 2025), we measure computational cost in FLOPs per iteration and memory cost in Bytes. Specifically, for each baseline, we report the relative FLOPs and relative Bytes in comparison to supervised fine-tuning (SFT), which only requires a single forward and backward pass without any extra computation and memory overhead.
Comparison of Computational Cost $\mathcal{C}$.
We first compare the computational cost of FL methods. PerAda (Xie et al., 2024) incurs approximately twice the computational cost compared to other baselines, since it sequentially updates both the personalized adapter and the local adapter. Similarly, FedDAT (Chen et al., 2024) independently optimizes the local adapter and the dual adapter teacher (DAT), which also results in a double computational cost. FedMKT, PerAda, and TAKFL perform knowledge distillation and transfer using logits (e.g., from public data), which introduces additional forward computation for logit extraction.
Figure 9: Relative Computational Cost to SFT vs. Number of clients of FedMosaic.
Our proposed FedMosaic incurs a minor additional computational cost due to (i) gradient computation from the frozen pre-trained model used for measuring task similarity and (ii) dual adapter structure, but this overhead amounts to only approximately 9.8% more FLOPs. The reason for the small increase in computation is that (i) we apply 60% gradient compression for similarity calculation in RELA, (ii) we only use the last layer gradient of the small-scale pre-trained frozen model, and (iii) we perform gradient computation once every 10 batch iterations rather than on all batches for computational efficiency, as mentioned in Sec. 4.1. Note that while FedMosaic also introduces a dual adapter structure, i.e., maintaining the local adapter and the global adapter separately, similar to PerAda, FedDAT, and FedDPA, we optimize the dual-adapter at once using a learnable balancing parameter in PQ-LoRA, which adaptively balances local and global output. In contrast, the baselines separately optimize each adapter, thus incurring a double computational cost. Thanks to these designs, the relative overhead remains approximately at 10% even under large-scale client setups (e.g., 1000 clients), as shown by the blue line in Fig. 9. Finally, the cost can be further reduced by adopting smaller or quantized $W_{s}$, as discussed in Sec. A.20. For example, using a LLaVA-Qwen-0.5B-4bit model as $W_{s}$ reduces the overhead to approximately 5% with 1000 clients (yellow line in Fig. 9), while achieving performance comparable to LLaVA-Llama-1B (Sec. A.20).
Comparison of Memory Cost $\mathcal{M}$.
We then compare the memory cost of FL methods. FedMKT incurs a marginal additional memory overhead, as it requires storing logits from clients for knowledge aggregation. DITTO, FedSim, PerAda, FedDAT, FedDPA, as well as FedMosaic employ a dual adapter structure, which maintains both local and global adapters separately, to preserve global knowledge. It incurs additional memory cost, but since we adopt the LoRA adapter (Hu et al., 2022), which occupies significantly less memory compared to pre-trained weights (Qi et al., 2024), the dual adapter only consumes about 5% additional memory cost compared to using a single LoRA adapter.
A.9 Details of Communication Costs in FedMosaic
The only additional transmission in FedMosaic compared to Vanilla in Tab. 7 (i.e., FedMosaic w/o RELA and w/o PQ-LoRA) is a single EMA-updated gradient vector per client, adding 8.6% overhead compared to sending only local LoRA parameters. However, as mentioned in Sec. 4.1, we apply gradient compression by randomly selecting only $N_{s}\%$ of the client-specific gradient vectors, thereby reducing communication costs. With $N_{s}=40\%$, the communication overhead drops to 3.4%.
Figure 10: Communication Cost vws. Number of clients of FedMosaic.
Moreover, while baselines transmit full LoRA modules (i.e., $A\in\mathbb{R}^{r\times d}$ and $B\in\mathbb{R}^{d\times r}$) across all layers, FedMosaic freezes $A$ and $B$ in PQ-LoRA layers and transmits only $P\in\mathbb{R}^{r\times r}$ and $Q\in\mathbb{R}^{r}$, significantly reducing 14.3% communication cost. Combining the marginal overhead from client-specific vectors with the reduction from PQ-LoRA, FedMosaic consequently achieves 10.9% lower communication cost than even the most efficient baseline (i.e., FedAvg), regardless of the number of clients, ensuring scalability and communication efficiency (Blue line in Fig. 10).
A.10 Empirical Analysis of Block-Wise Aggregation
@@
To identify layer-wise correspondences between depth-heterogeneous models, we analyze representation alignment using CKA (Kornblith et al., 2019). Specifically, we measure similarity across layers within the Llama-3 family (1B, 3B, 8B) and the Qwen-2.5 family (0.5B, 1.5B, 3B), as illustrated in Fig. 11. As shown in the figure, layers with the same relative depth exhibit strong alignment, indicating approximately linear alignment within both the Llama-3 and Qwen-2.5 families. Moreover, we observe near-linear alignment even across families, i.e., between Llama-3 and Qwen-2.5, despite weaker linearity than intra-family alignment. This empirical analysis supports our block-wise aggregation of PQ-LoRA. We provide an illustration of the block-wise PQ-LoRA in Fig. 12.
@@
A.11 Experimental Results in More Diverse Heterogeneous PFL Scenarios
@@
DRAKE-Dynamic DRAKE-Static
@@
Method $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]
@@
A.12 Experiment Results on Homogeneous PFL Setup
In addition to the heterogeneous PFL scenario (Sec. 6.2), where both data distributions and model architectures vary across clients, we also evaluate FedMosaic in a homogeneous PFL setting, where clients share the same model architecture but have heterogeneous data distributions. We assume all clients use LLaVA-3B models and summarize the results in Tab.17. Consistent with results in the heterogeneous setup, FedMosaic significantly outperforms baselines in generalization ability (i.e., ‘Others’) and even surpasses SFT in personalization performance (i.e., ‘Self’).
DRAKE-Dynamic HFLB-Dynamic
@@ A.12 Experiment Results on Homogeneous PFL Setup
Method $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]
DRAKE-Static HFLB-Static
@@
Method $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]
@@
A.13 Experiment Results on Qwen-based LLaVA
@@
As shown in the table, FedMosaic significantly outperforms baselines in both homogeneous and heterogeneous PFL scenarios, consistent with results from LLaVA-Llama3 variants. This demonstrates that our proposed PQ-LoRA consistently facilitates knowledge sharing across heterogeneous models and RELA reduces interference during local model aggregation, regardless of architecture.
DRAKE-Homonegeous DRAKE-Heterogeneous
@@
Method $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]
@@
A.14 Experiment Results in Cross-Family Heterogeneous PFL Scenarios
In addition to Tab. 4, we further evaluate FedMosaic under cross-family heterogeneity using clients with either Qwen- and Llama-based LLaVAs on DRAKE-Dynamic. As shown in Tab. 19, FedMosaic consistently outperforms all baselines, demonstrating its generalizability and transferability beyond the same family heterogeneous models (e.g., LLaVA-Llama-1B, LLaVA-Llama-3B).
Llama 3B / Qwen 1.5B Llama 1B / Llama 3B / Qwen 1.5B
Self Others Self Others
Method $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]
Table 19: Quantitative comparison in cross-family heterogeneous PFL on DRAKE-dynamic. ‘Self’ denotes evaluation on a client’s own data, while ’Others’ denotes evaluation on data from other clients. ‘Llama 3B / Qwen 1.5B’ experiment is with 6 clients using LLaVA-Llama3.2-3B model and 4 clients using LLaVA-Qwen2.5-1.5B model, while ‘Llama 1B / Llama 3B / Qwen 1.5B’ experiment is with 2 clients using LLaVA-Llama3.2-1B model, 5 clients using LLaVA-Llama3.2-3B model, and 3 clients using LLaVA-Qwen2.5-1.5B model. SFT refers to supervised fine-tuning on each client’s data without cross-client knowledge sharing.
A.15 Experimental Results on Text-Only Benchmarks
In addtion to Fed-LLM-Large (Sec.6.2), we evaluate FedMosaic on other text-only PFL benchmarks, Fed-Aya and Fed-Scope, and summarize the results in Tab. 20. As shown in the table, FedMosaic outperforms baselines in both personalization and generalization, consistent with results from other MLLM PFL benchmarks, such as DRAKE and HFLB.
Fed-Aya Fed-Scope
Self Others Self Others
Method $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]
Table 20: Quantitative comparison of heterogeneous LLM clients on the text-only benchmarks. In Fed-aya experiment, 4 clients use Llama-3.2-1B and 4 clients use Llama-3.2-3B. In Fed-Scope experiment, 3 clients use Llama-3.2-3B and 2 clients use Llama-3.1-8B.
A.16 Client-Wise Accuracy